OpenStudy (anonymous):

Could some one explain how to solve this? ''on the same set of axes, sketch y=x^2 and the function: y=-1/-2 x^2 And Y=(x+2)^2 +5 We just got this in math class and we don't have a math teacher, no one understands this since they don't mention how to do this in the book....

OpenStudy (mani_jha):

Can you sketch y=x^2?

OpenStudy (mani_jha):

The graph of y=(x+2)^2 + 5 will be a parabola at the vertex (-2,5)(just like the vertex of y=x^2 is (0,0)). Can you draw that?

OpenStudy (anonymous):

I see it now.... ^^